### 02 Feb 2024 v1

# HimiRec: Modeling Hierarchical Multi-interest for Recommendation

Abstract.
Industrial recommender systems usually consist of the retrieval stage and the ranking stage, to handle the billion-scale of users and items. The retrieval stage retrieves candidate items relevant to user interests for recommendations and has attracted much attention. Frequently, users show hierarchical multi-interests reflected in a heavy user of a certain NBA team Golden State Warriors in Sports, who is also a light user of almost the whole Animation. Both Sports and Animation are at the same level. However, most existing methods implicitly learn this hierarchical difference, making more fine-grained interest information to be averaged and limiting detailed understanding of the user’s different needs in heavy interests and other light interests. Therefore, we propose a novel two-stage approach to explicitly modeling hierarchical multi-interest for recommendation in this work. In the first hierarchical multi-interest mining stage, the hierarchical clustering and transformer-based model adaptively generate circles or sub-circles that users are interested in. In the second stage, the partition of retrieval space allows the EBR models to only deal with items within each circle and accurately capture users’ refined interests. Experimental results show that the proposed approach achieves state-of-the-art performance. Our framework has also successfully deployed at Lofter (one of the largest derivative content communities with 10 million monthly active users) for over four months.
Recommender Systems, Multi-interest Recommendation, Generative Retrieval, Embedding-based Retrieval
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The Web Conference; May 13 - May 17, 2023; Singaporeprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Figure 1. The tree of circle “Animation” with sub-circles and an content flow example in “Animation, Japanese Animation, Detective Conan, Conan”. Note that each circle (including sub-circle) has a content flow and the content of the upper layer circle is a mixture of the contents of sub-circles.
Lofter, one of the largest derivative content communities, has over 10 million monthly active users who interact with more than one billion posts. Circles with different sub-circles were derived spontaneously to share various interests by users. Take Lofter as an example, Figure 1 shows circle “Animation” with sub-circles, for discussion of interest in different anime from different countries. Frequently, users show refined interests in different circles. For example, a heavy user of a certain NBA team Golden State Warriors (usually familiar with the slang in the sub-circle “Golden State Warriors” of Sports circle), is also a light user of almost the whole Animation circle. Both Sports and Animation are the same level circle. We call it hierarchical multi-interest, requiring personalized modeling of a user’s heavy interests and other light interests.
Approaches for multi-interest recommendation provide a feasible solution. Recently, most prior work has inferred multiple high dimensional embeddings to represent a user’s interests (Cen et al., 2020; Pal et al., 2020; Li et al., 2023; Xie et al., 2023) to support retrieving items from the entire library, overcoming the bottleneck of using a single generic user embedding. Others explicitly partition the library and retrieve items at all sub-libraries with embeddings produced by the same user sequence to extend multi-interest recommendation (Zhang et al., 2023). However, their approaches suffer from weakness: implicitly learn this hierarchical difference, making more fine-grained interest information to be averaged and limiting detailed understanding of the user’s different needs in heavy interests and other light interests.
To address the issues mentioned above, we propose HimiRec, a novel two-stage retrieval method for hierarchical multi-interest recommendation. In the first hierarchical multi-interest mining stage, an interest tree for all items was constructed by hierarchical clustering. Each item obtains a Semantic Category ID to model the cycle or sub-circle it belongs to. After that, we build a transformer-based model with early stopping to directly generate Semantic Category IDs that the user may be interested in based on user behavior to achieve hierarchical multi-interest mining. In the second items retrieval stage, the library is explicitly partitioned into sub-libraries based on the item categories tree. For each user, we capture the items the user consumed that fall into the categories generated in the first stage to build multiple embeddings, to perform ANN searches in corresponding sub-libraries. This allows longer behavior sequences in each category for more refined and personalized recommendations.
Our main contributions are as follows:
  • A novel two-stage retrieval model for hierarchical multi-interest recommendation was proposed.
  • The hierarchical clustering and transformer-based model adaptively generate circles or sub-circles that users are interested in, contributing to refined multi-interest mining.
  • Our approach achieves state-of-the-art performance, excelling in both offline datasets and real-world online A/B testing scenarios. Provide industrial solutions for refined hierarchical multi-interest recommendations in similar scenarios.
2. Method
2.1. Overall Architecture
Suppose there are a set of users $\mathcal{U}$ and a set of items $\mathcal{I}$. Denote the user-side features as $\bm{x}_{u}$ for each user $u\in\mathcal{U}$ and item-side features as $\bm{x}_{i}$ for each item $i\in\mathcal{I}$. Embedding-based retrieval (EBR) methods use a user encoder $f$ and an item encoder $g$ to transform user and item features into embeddings $\bm{e}_{u}=f(\bm{x}_{u})$ and $\bm{e}_{i}=g(\bm{x}_{i})$ respectively, where $i\in[1:\lvert\mathcal{I}\rvert]$. The distance (most commonly inner product) between user embedding $\bm{e}_{u}$ and item embedding $\bm{e}_{i}$ is used to capture the relevance of item $i$ to user $u$ denoted by $r_{ui}$.
As shown in Figure 2 and Figure 3, our proposed method consists of two components: (1) The multi-interest mining stage achieves multi-interest mining for users by a generated-based model; (2) The item retrieval stage performs ANN searches in partitioned sub-libraries to achieve refined and personalized recommendations.
2.2. Hierarchical Multi-interest Mining
Semantic Category ID Assignment. Semantic Category ID was assigned for each item based on the item features (e.g. titles, descriptions), representing the circle the item belongs to. Specifically, mapping the item features to embedding vectors and then clustering via the hierarchical $k$-means algorithm. As shown in Algorithm 1, all items are first clustered into $k$ clusters. For clusters with more than $c$ items, the $k$-means algorithm is applied recursively. Each cluster in the $i$-th layer is assigned a number $r_{i}$ starting from 0 to at most $k$-1. In this way, we organize all items into a tree structure, and the semantic category ID set $L$ of the leaf circles is obtained, where the category of each leaf circle in $L$ is represented by path $l=\{r_{1},\dots,r_{m}\}$. Futhermore, each item in the leaf cluster $r_{m}$ in $l$ is assigned any path $l^{\prime}=\{r_{1},\dots,r_{p}\}$, where $1\leq p\leq m,p\in\mathbb{N}^{+}$, as the semantic category ID.
Figure 2. The architecture of the multi-interest mining stage of the proposed approach, HimiRec.
Figure 3. The architecture of the item retrieval stage of the proposed approach, HimiRec.
0:     Item embedding $\bm{e}_{1:\lvert\mathcal{I}\rvert}$umber of clusters $k$Recursion terminal condition $c$ Prefix $l_{prefix}$ (Initialized to an empty string)
0:    Semantic category ID $L$ of leaf circles
0:    GenerateSemanticCategoryID($\bm{e}_{1:\lvert\mathcal{I}\rvert}$, $l_{prefix}$)
  $C_{1:k}$leftarrow $KMeansCluster(\bm{e}_{1:\lvert\mathcal{I}\rvert},k)$
  $L$leftarrow $\varnothing$
  for i $\in$ [0, $k-1$do
     $r_{i}$leftarrow $i$
     if $|C_{i+1}|$ $c$ then
        $l$leftarrow GenerateSemanticCategoryID$(C_{i+1},l_{prefix}+str(i))$
     else
        $l$leftarrow ConcatString($l_{prefix},r_{i}$)
     end if
     $L$leftarrow $L$.Append($l$)
  end for
  return  $L$
Algorithm 1 Hierarchical $k$-means.
Multi-interest Generation. A Transformer-based sequence-to-sequence model is trained on the user behavior sequences $\bm{s}^{u}$ to predict the possible circles that the user is interested in by generating corresponding semantic category IDs. Note that $\bm{s}^{u}$, including the sampled long-term sequence and all short-term sequence items, is obtained from the user’s lifelong behaviors $\mathcal{I}^{u}$. The generative-based model can adaptively learn the users’ refined interest reflected in generating IDs of different lengths via an early stopping mechanism (that generates the end token). Notice that we did not share the embedding space of the encoder and decoder since the vocabulary spaces of the user’s interaction history and semantic category IDs are quite different. The implementation details are as follows.
Given a user behavior sequence $\bm{s}^{u}$, the probability of generating a semantic category ID $l^{\prime}$ is calculated as follows:
(1) $h=\text{TransformerEncoder}(\bm{s}^{u};\theta_{encoder})$
(2) $P(l^{\prime}|\bm{s}_{u},\theta)=\prod_{i=1}^{p}{P(r_{i}|h,r_{1},r_{2},...,r_{i-1},\theta_{decoder})}$
where $h$ is the representation output from encoder; $\theta$ denotes the whole model parameters. In the training stage, $p=m$ represents the semantic category ID of the leaf circle to which the items that the user $u$ has interacted historically belong. In the inference stage, we perform a beam search on the decoder network with a prefix tree for valid identifiers, to obtain the user’s refined interest $L_{u}=\{l^{\prime}_{1},\dots,l^{\prime}_{\lvert\text{beam size}\rvert}\}$, where $1\leq p\leq m,p\in\mathbb{N}^{+}$.
In particular, the same number appearing at different positions in a semantic category ID is different. For instance, as shown in Figure 2, the token “$t_{(2,4)}$” and “$t_{(3,4)}$” are different tokens in the vocabulary space, corresponding to the circle IDs in layer2 and layer3, respectively. Together (with token “$t_{(1,2)}$”) they represent the the semantic category ID “$\{2,4,4\}$”.
2.3. Refined Multi-interest Retrieval
A modified two-tower model was used for multi-interest retrieval. First partition the whole library $\mathcal{I}$ into sub-libraries $\{\mathcal{I}_{1},\dots,\mathcal{I}_{\lvert L\rvert}\}$ based on the semantic category ID set of the leaf circles $L$ defined in Section 2.2. Note that sub-library $\mathcal{I}_{i}$ can be merged at the upper layer flexibly based on $L_{u}$ generated in the first stage, and $\{\mathcal{I}^{1},\dots,\mathcal{I}^{\lvert L_{u}\rvert}\}$ are obtained. Secondly, the user’s lifelong behavior sequences $\mathcal{I}^{u}$ is split following the Key-Key-Value data structure: the first key is the user ID, the second keys are semantic category IDs $L_{u}$ generated in the first stage and the last values are the user behavior items $\mathcal{I}^{u}_{l^{\prime}_{n}}$ that fall into each semantic category ID $l^{\prime}_{n}$. This allows a longer behavior sequence for each category compared to the previous method and massive noise existing in the lifelong user behavior that hinders refined interest modeling can be filtered by the split strategy. After that, for each Key-Key-Value, with a separate sequence, a more refined user embedding is learned, and further a user has multiple distinction user embeddings. Then retrieve relevant items in the sub-library $\mathcal{I}^{i}$ that each user embedding corresponds to instead of the whole library. Lastly, the beam probability in the first stage will be used for assigning quotas to merging results. We only use items from the same category of positive items as negative items (that are mostly hard negatives) for training as well. The training loss is
(3) $\mathcal{L}=-\sum_{u\in\mathcal{U}}\sum_{n=1}^{\lvert L_{u}\rvert}\sum_{i^{+}\in\mathcal{I}^{u}\cap\mathcal{I}^{n}}\Bigl{[}\log(\sigma(r_{ui^{+}}))+\mathop{\mathbb{E}}_{i^{-}\sim\mathcal{I}^{n}\setminus\mathcal{I}^{u}}\bigl{[}\log(1-\sigma(r_{ui^{-}}))\bigr{]}\Bigr{]}.$
where $\mathcal{I}^{n}$ represents the sub-library of the semantic category ID $l^{\prime}_{n}$. During employment, item embeddings $\{\bm{e}_{i}\}_{i\in\mathcal{I}}$ are calculated beforehand in an offline or near-line system and indexed by approximate nearest neighbor (ANN) search system (e.g. FAISS (Johnson et al., 2021)). Therefore, only the user encoder $f(\mathcal{I}^{u}_{l^{\prime}_{n}};\bm{x}_{u})$ needs to be evaluated in real-time, and the ANN system can efficiently retrieve the nearest items in sub-linear time.
The partition of item space allows deal only with items within each category and accurately captures users’ refined interests in the target category, meeting users’ different needs in different circles.
3. Offline Experiments
Datasets. We conduct extensive experiments on two public datasets. Table 2 summarizes the detailed statistics of the datasets.
  • MovieLens-1M (Harper and Konstan, 2016) is a public movie rating dataset commonly used in recommender systems. Randomly sample 80% of the users as the training set and 20% of the users as the test set. Learn only from the user sequences without user ID.
  • LofterData collects user behaviors from the logs of Lofter for 15 days with 12,000 users (that clicks and daily interactions are less than $20$ and $5$ respectively were filtered out). The 15th-day data is retained for testing.
Parameter settings. We set $k$ as 4 and 6, c as 250 and 100,000 for ML-1M and real Lofter scenery respectively. During inference, each method generates a candidate set of $M$ relevant items and we evaluate its Recall@$M$ and HitRate@$M$ on the following $N$ disjoint behaviors. We set $M\in\{20,50\}$, $N=10$ for ML-1M and $M\in\{500,1000\}$, $N=25.3$ in average for LofterData.
Baselines. We compare our approach with the common baselines used in the industry: MIND (Li et al., 2019) uses capsule networks with the routing mechanism to group user behaviors into multiple clusters and obtain multiple embeddings for retrieval. ComiRec-SA (Cen et al., 2020) uses multi-head attention mechanisms to generate multiple embeddings to capture their diverse interests. Divide&Conquer (Zhang et al., 2023) divides the whole set of candidate items into multiple clusters and runs EBR to retrieve relevant candidates from each cluster in parallel.
Table 1. Performance comparison on MovieLens-1M and LofterData. All the numbers are percentage numbers with ‘%’ omitted.
Method MovieLens-1M LofterData
Recall@20 HitRate@20 Recall@50 HitRate@50 Recall@500 HitRate@500 Recall@1000 HitRate@1000
MIND 4.557 23.372 11.605 43.294 24.754 61.158 38.719 71.660
ComiRec 7.744 44.238 17.165 65.820 26.010 61.467 40.410 70.156
Divide&Conquer 7.107 48.014 15.465 69.889 27.722 68.035 44.965 75.95
HimiRec-2beams 9.119 50.878 19.263 73.542 32.428 66.332 46.581 72.781
HimiRec-4beams 9.401 51.975 20.498 76.176 37.083 66.561 59.376 73.667
HimiRec-6beams 9.461 52.571 20.630 76.207 39.978 67.740 65.430 73.748
Improvement +17.8% +15.0% +12.2% +11.7% +24.7% +7.9% +15.3% +3.7%
Main Result. Table 1 summarizes the performance of the proposed approach as well as baselines on MovieLens-1M and LofterData. First, the proposed approach outperforms all the baseline methods on Recall consistently on both datasets and achieves up to $17.8\%$ and $24.7\%$ improvements over the baseline model ComiRec on MovieLens-1M and LofterData datasets, respectively. However, our HitRate is not as good as Divide&Conquer. The reason could be that Divide&Conquer searched in almost all sub-libraries. However, our method achieves better performance under the trade-off in Recall and HitRate. Secondly, as the number of interest embeddings increases, our model performance becomes better. An Ablation Study was done to assess the proposed HimiRec approach. Results in table 3 show that by performing ANN search in the whole library rather than in the corresponding sub-libraries, the Recall@500 and Recall@1000 performance drops by up to $-8.0\%$ and $-4.5\%$ respectively on LofterData. What’s more, results show that performing ANN search with user embedding produced by a unified user sequence rather than the sequence under the corresponding category, the overall performance drops by up to $-4.4\%$ and $-2.3\%$ on LofterData, respectively. The reason why the improvement is not significant could be that the sequence length in specific categories on the offline data set is shorter and the features are sparser.
Table 2. Statistics of datasets (after preprocessing).
Dataset #Users #Items #Interactions #Leaf Circles
ML-1M 6,040 3,411 892,458 28
LofterData 12,000 798,949 808,619 176
Table 3. Ablation study on the Lofterdata (with ‘%’ omitted).
Method Recall@500 Recall@1000
w/o sub-libraries (2beams) 30.028 44.590
w/o behavior splitting (2beams) 31.061 45.534
4. Online A/B Experiments
The proposed approach has also been tested in a strict online A/B experiment at Lofter for a month (from 2023-08-03 to 2023-09-03), and involved over 10% users in the experiment group, yielding statistically significant at $p<0.05$ for all experimental results. In the experiment group, the proposed retrieval method serves as one of the candidate sources in the candidate generation phase. Compared with the control group (that ComiRec was deployed), our method leads to significant improvements on most of the key user engagement metrics as shown in Table 4. Currently, HimiRec has been deployed online and serves the main scene daily. Figure 4 shows an example result of the multi-interest mining stage. The model can capture the user’s hierarchical multi-interests.
Table 4. Results of the online A/B experiment. Note that: 1) Diversity indicates the Ratio-based diversity (Wang et al., 2022) with its numerator replaced by the number of unique items the user consumed, measuring the diversity and broadness of the contents that a user consumed. 2) User Activity records the frequency of users visiting the app in a week.
User Activity Click Engagement Retention Diversity
+0.31% +0.76% +2.00% +0.45% +1.26%
Figure 4. (a) Part of interaction sequence of a certain user in Lofter; (b) The Semantic Category IDs generated by the Multi-interest Generation stage and the items retrieved by the Refined Multi-interest Retrieval stage.
5. Conclusion
This paper proposes a production-proven two-stage retrieval model for hierarchical multi-interest recommendation. The hierarchical clustering and transformer-based model adaptively generate circles or sub-circles that users are interested in. The partition of retrieval space allows the EBR models to only deal with items within each category and accurately capture users’ refined interests. Extensive results in both offline and online A/B experiments demonstrate the effectiveness of the proposed solution.